# Generalized and Efficient 2D Gaussian Splatting for Arbitrary-scale Super-Resolution

### v2

@@ 5 Conclusion and Limitation
We presented GSASR, a brandly-new 2D GS-based ASR model. To adapt GS to the ASR task, we firstly elaborately designed an architecture to covert an LR input to image-conditioned Gaussians, then implemented an efficient differentiable scale-aware 2D GPU/CUDA-based rasterization to render images with both high quality and fast speed. Through sampling values from Gaussians, we could render an output image with arbitrary magnification factors. Extensive experiments demonstrated that our GSASR model has a much more powerful representation capability, together with more friendly computational costs, than Implicit Neural Function based methods.
Limitations. GSASR still has some limitations. First, its performance depends on the number of Gaussians, especially under ultra-high scaling factors. While more Gaussians could bring better performance, it will sacrifice the inference speed. Second, to convert an image into Gaussians, a large number of parameters are used to fit the complex mapping function. How to reduce the number of parameters needs further investigation.

### v3

@@ 5 Conclusion and Limitation
We presented GSASR, a brandly-new 2D GS-based ASR model. To adapt GS to the ASR task, we first elaborately designed an architecture to covert an LR input to image-conditioned Gaussians, then implemented an efficient differentiable scale-aware 2D GPU/CUDA-based rasterization to render images with both high quality and fast speed. Through sampling values from Gaussians, we could render an output image with arbitrary magnification factors. Extensive experiments demonstrated that our GSASR model has much more powerful representation capability, together with more friendly computational costs, than implicit neural function based methods.
Limitations. GSASR still has some limitations. First, its performance depends on the number of Gaussians, especially under ultra-high scaling factors. While employing more Gaussians could bring better performance, it will sacrifice inference speed. Second, to convert an image into Gaussians, a large number of parameters are used to fit the complex mapping function. How to reduce the number of parameters needs further investigation.